OpenStudy (anonymous):

I got a question I know how to do point slope formula with two coordinates but if im rearranging that formula in a standard form how would i get rid of a slope thats a fraction

OpenStudy (anonymous):

cross multiplication

OpenStudy (anonymous):

cross multiply with what though I mean my best guess of coss multiplying would be to multipy everything in equation by denominator except for anything in parenthesis

OpenStudy (anonymous):

and get rid of fractthat way

OpenStudy (anonymous):

and get rid of fraction that way

OpenStudy (anonymous):

example y - 3 = 1/3(x - 4) now would i multiply everything by denominator 3y - 9 = 1(x - 4)

OpenStudy (anonymous):

to get rid of slope fraction

OpenStudy (anonymous):

yes..